OpenStudy (anonymous):

Find a pattern for each sequence. Use the pattern to show the next two terms. 3. -3, 6, -12, 24, -48, … 4. 1, 5, 30, 210, 1680, …

OpenStudy (tkhunny):

You can often track this sort of thing down with successive differences. -3 - 3 = -6 = -2*3 6 - (-3) = 9 = +3*3 -12 - 6 = -18 = -6*3 = -2*3*3 <== Multiplied previous by -2 24 - (-12) = 36 = +12*3 = 2*2*3*3 <== Multiplied previous by -3 -48 - 24 = -72 = -24*3 = -2*2*2*3*3 <== Multiplied previous by -2 Maybe the differences are not very helpful on this one. You do just have to play around with it. 3 -3 = -1(3) 6 = -2*(-3) -12 = -2(6) 24 = -2(-12) -48 = -2(24) Again, except for the 1st one, we're there. Just toy with it. You'll think of something. Occasionally, there will be two sequences interwoven. 3, 6, 24 ==> 3, 3*2, 3*8 ==> 3*2^0, 3*2^1, 3*2^3, maybe 3*2^6 is next? -3, -12, -48 ==> -3, -3*4. -3*16, maybe -3*64 is next?

OpenStudy (anonymous):

you have to do all the steps with diffenrt numbers